# A congruence subgroup property for symmetric mapping class groups

### v1

@@ 1. Introduction
Let $S$ be an oriented connected surface of negative Euler characteristic and finite type and let $G$ be a finite group acting faithfully on $S$. The group $G$ then identifies with a finite subgroup of the mapping class group ${\Gamma}(S)$ of the surface $S$ and let us denote by ${\Gamma}_{G}(S)$ the centralizer of $G$ in ${\Gamma}(S)$. Let $\Pi$ be the fundamental group of $S$ for some choice of base point. There is a natural faithful representation $\rho_{S}\colon\thinspace{\Gamma}(S)\hookrightarrow\operatorname{Out}(\Pi)$ which restricts to a faithful representation $\rho_{S,G}\colon\thinspace{\Gamma}_{G}(S)\hookrightarrow\operatorname{Out}(\Pi)$.
@@
(1) $\hat{\rho}_{S}\colon\thinspace{\widehat{\Gamma}}(S)\to\operatorname{Out}({\widehat{\Pi}})\hskip 28.45274pt\mbox{and}\hskip 28.45274pt\hat{\rho}_{S,G}\colon\thinspace{\widehat{\Gamma}}_{G}(S)\to\operatorname{Out}({\widehat{\Pi}}).$
The congruence subgroup problem for the mapping class group ${\Gamma}(S)$ (resp. for ${\Gamma}_{G}(S)$) asks whether the homomorphism $\hat{\rho}_{S}$ (resp. $\hat{\rho}_{S,G}$) is faithful. The first question is known to have a positive answer for $g(S)\leq 2$ (cf. [1], [2], [6]), where $g(S)$ denotes the genus of $S$.
@@
Let $S$ be an oriented connected hyperbolic surface of finite type and $G$ a finite group acting faithfully on $S$. If the congruence subgroup property holds for the mapping class group ${\Gamma}(S_{/G}^{\circ})$ of the surface $S_{/G}^{\circ}$, then it holds for the centralizer ${\Gamma}_{G}(S)$ of $G$ in the mapping class group ${\Gamma}(S)$ of $S$.
@@
If $g(S_{/G})\leq 2$, then the congruence subgroup property holds for ${\Gamma}_{G}(S)$.
For a subgroup $H$ of a group $G$, we denote by $Z_{H}(G)$ and $N_{H}(G)$, respectively, the centralizer and the normalizer of $H$ in $G$. For a subset $A$ of a profinite group $R$, we denote by $\overline{A}$ the closure of $A$ for the profinite topology of $R$. We then have:
@@
    The natural monomorphism ${\Gamma}(S,\partial S)\hookrightarrow{\widehat{\Gamma}}(S,\partial S)$ induces a bijection between conjugacy classes of solvable finite subgroups. In particular, ${\Gamma}(S,\partial S)$ is solvable finite subgroup conjugacy separable.
@@
    Torsion elements are conjugacy distinguished.
@@
    If, moreover, $G$ is solvable, we also have $N_{{\widehat{\Gamma}}(S,\partial S)}(G)=\overline{N_{{\Gamma}(S,\partial S)}(G)}$.
@@ 2.1. Normalizers and centralizers of finite subgroups of the mapping class group of a surface
(2) $1\to G\to N_{{\Gamma}(S)}(G)\to{\Gamma}(S_{/G}^{\circ})_{K}\to 1$
Let $Z(G)$ be the center of the group $G$ and let ${\Gamma}^{K}$ be the kernel of the natural representation ${\Gamma}(S_{/G}^{\circ})_{K}\to\operatorname{Out}(\left.\Pi_{/G}\right/K)$. The centralizer ${\Gamma}_{G}(S)\leq N_{{\Gamma}(S)}(G)$ of the subgroup $G$ in the mapping class group ${\Gamma}(S)$ is then described by the short exact sequence:
(3) $1\to Z(G)\to{\Gamma}_{G}(S)\to{\Gamma}^{K}\to 1$
@@
$N_{{\Gamma}(S^{\circ})}(G)\cong N_{{\Gamma}(S)}(G)\hskip 28.45274pt\mbox{and}\hskip 28.45274pt{\Gamma}_{G}(S^{\circ})\cong{\Gamma}_{G}(S).$
@@ 2.2. The congruence topology
The congruence topology on ${\Gamma}(S)$ is the subspace topology induced by the standard profinite topology on $\operatorname{Out}({\widehat{\Pi}})$ via the natural monomorphism ${\Gamma}(S)\hookrightarrow\operatorname{Out}({\widehat{\Pi}})$. The congruence topology on ${\Gamma}(S)$ then induces a topology on its subgroups which we also call the congruence topology.
Let ${\Gamma}_{G}^{(m)}(S)$ be the kernel of the natural representation ${\Gamma}_{G}(S)\to\operatorname{GL}(H_{1}(S,\mathbb{Z}/m))$ and let ${\Gamma}^{K,(m)}$ be its image via the natural epimorphism ${\Gamma}_{G}(S)\to{\Gamma}^{K}$. For $m\geq 3$, we have ${\Gamma}_{G}^{(m)}(S)\cap G=\{1\}$ and so the map ${\Gamma}_{G}(S)\to{\Gamma}^{K}$ induces an isomorphism ${\Gamma}_{G}^{(m)}(S)\cong{\Gamma}^{K,(m)}$.
@@
For all $m\geq 2$, the finite index subgroups ${\Gamma}_{G}^{(m)}(S)$ of ${\Gamma}_{G}(S)$ and ${\Gamma}^{K,(m)}$ of ${\Gamma}(S_{/G}^{\circ})$ are both open for the respective congruence topologies.
@@
The statement is obvious for ${\Gamma}_{G}^{(m)}(S)$. For ${\Gamma}^{K,(m)}$, it can be proved with the same argument of [3, Corollary 2.3]. ∎
@@ 2.3. The centralizer of a finite subgroup in the profinite completion of a good group
Let $G$ be a good, residually finite group of finite virtual cohomological type and let $\iota\colon\thinspace G\hookrightarrow\widehat{G}$ be the natural monomorphism. Then, for every finite subgroup $H$ of $G$, we have the identity $C_{\widehat{G}}(\iota(H))=\overline{\iota(C_{G}(H))}$.
@@
By [5, (iii) of Theorem A], for all primes $p>0$ and every $p$-torsion element $x$ of $G$, there holds $C_{\widehat{G}}(\iota(x))=\overline{\iota(C_{G}(x))}$. The finite group $H$ is generated by a finite subset $\{x_{1},\dots,x_{k}\}$ of $p$-torsion elements (for, possibly, different primes $p>0$). We then have:
$C_{\widehat{G}}(\iota(H))=\bigcap_{i=1}^{k}C_{\widehat{G}}(\iota(x_{i}))=\bigcap_{i=1}^{k}\overline{\iota(C_{G}(x_{i}))}=\overline{\iota(C_{G}(H))}.$
@@ 3.1. The case of a free action
The congruence subgroup property for ${\Gamma}_{G}(S)$ is equivalent to the statement that the congruence topology on this group, as defined in Section 2.2, is equivalent to the profinite topology. Since, by Lemma 2.2, the subgroup ${\Gamma}_{G}^{(m)}(S)$ of ${\Gamma}_{G}(S)$ is open for the congruence topology, the congruence subgroup property for ${\Gamma}_{G}(S)$ follows if we show that the homomorphism:
$\hat{\rho}_{S,G}^{(m)}\colon\thinspace{\widehat{\Gamma}}_{G}^{(m)}(S)\to\operatorname{Out}({\widehat{\Pi}}),$
induced by the restriction of $\rho_{S,G}$ to ${\Gamma}_{G}^{(m)}(S)$, is injective for some $m\geq 2$.
Since, by hypothesis, the representation $\hat{\rho}_{S_{/G}}\colon\thinspace{\widehat{\Gamma}}(S_{/G})\to\operatorname{Out}({\widehat{\Pi}}_{/G})$ is faithful, its restriction $\hat{\rho}_{S_{/G}}^{(m)}\colon\thinspace{\widehat{\Gamma}}^{K,(m)}\to\operatorname{Out}({\widehat{\Pi}}_{/G})$ is faithful as well for all $m\geq 2$.
For $m\geq 3$, the natural isomorphism ${\Gamma}_{G}^{(m)}(S)\cong{\Gamma}^{K,(m)}$ induces an isomorphism of profinite completions ${\widehat{\Gamma}}_{G}^{(m)}(S)\cong{\widehat{\Gamma}}^{K,(m)}$ and, since ${\widehat{\Pi}}\cong\widehat{K}$ identifies with a normal finite index subgroup of ${\widehat{\Pi}}_{/G}$ and the representations $\hat{\rho}_{S,G}^{(m)}$ and $\hat{\rho}_{S_{/G}}^{(m)}$ are compatible via these isomorphisms, the case of a free action of Theorem A follows from the well known lemma:
@@
Let ${\widehat{\Pi}}$ be a nonabelian profinite surface group, $N$ an open normal subgroup of ${\widehat{\Pi}}$ and $f\in\operatorname{Aut}({\widehat{\Pi}})$. If $f$ restricts to the identity on $N$, then $f$ is the identity.
@@ 3.2. The general case
The completions of the groups ${\Gamma}(S)$ and ${\Gamma}_{G}(S)$ with respect to the congruence topologies which we introduced in Section 2.2 are called procongruence completions and denoted by ${\check{\Gamma}}(S)$ and ${\check{\Gamma}}_{G}(S)$, respectively. The group ${\check{\Gamma}}(S)$ is also called the procongruence mapping class group. The congruence subgroup problems for ${\Gamma}(S)$ and ${\Gamma}_{G}(S)$ can then be formulated as the questions of whether the natural epimorphisms ${\widehat{\Gamma}}(S)\to{\check{\Gamma}}(S)$ and ${\widehat{\Gamma}}_{G}(S)\to{\check{\Gamma}}_{G}(S)$ are, respectively, isomorphisms.
By the results of Section 3.1 and with the notations of Section 1, we know that there is a natural isomorphism ${\widehat{\Gamma}}_{G}(S^{\circ})\cong{\check{\Gamma}}_{G}(S^{\circ})$. Therefore, in order to prove the general case of Theorem A, it is enough to show that the natural isomorphism $q^{\circ}\colon\thinspace{\Gamma}_{G}(S^{\circ})\stackrel{{\scriptstyle\sim}}{{\to}}{\Gamma}_{G}(S)$, induced by the natural epimorphism $q^{\circ}\colon\thinspace{\Gamma}(S^{\circ})\to{\Gamma}(S)$ (cf. Proposition 2.1), induces an isomorphism between the procongruence completions: $\check{q}^{\circ}\colon\thinspace{\check{\Gamma}}_{G}(S^{\circ})\stackrel{{\scriptstyle\sim}}{{\to}}{\check{\Gamma}}_{G}(S)$.
The set $R$ of ramification points of the quotient map $p_{G}\colon\thinspace S\to S_{/G}$ determines a point, which we also denote by $R$, on the unordered configuration space $S[r]$ of $r$ distinct points on the surface $S$, where $r:=\sharp R$.
Let $\Pi[r]:=\pi_{1}(S[r],R)$. From [6, Corollary 4.7 and Proposition 4.9], it then follows that there is a short exact sequence of profinite groups:
(4) $1\to{\widehat{\Pi}}[r]\to{\check{\Gamma}}(S^{\circ})\stackrel{{\scriptstyle\check{q}^{\circ}}}{{\to}}{\check{\Gamma}}(S)\to 1$
In order to prove Theorem A, it is then enough to show that the subgroup ${\check{\Gamma}}_{G}(S^{\circ})$ of ${\check{\Gamma}}(S^{\circ})$ has trivial intersection with the image of ${\widehat{\Pi}}[r]$ in ${\check{\Gamma}}(S^{\circ})$. Since ${\check{\Gamma}}_{G}(S^{\circ})$ is contained in the centralizer $Z_{{\check{\Gamma}}(S^{\circ})}(G)$, the conclusion follows if we prove that $Z_{{\check{\Gamma}}(S^{\circ})}(G)\cap{\widehat{\Pi}}[r]=\{1\}$.
@@
Proposition 2.1 implies that ${\Gamma}_{G}(S^{\circ})\cap\Pi[r]=\{1\}$ and then that $Z_{\Pi[r]\cdot G}(G)=Z(G)$. Since $\Pi[r]\cdot G$ contains, as a finite index subgroup, the fundamental group of the ordered configuration space of $n$ points on the surface $S$, which is a good, residually finite group of finite virtual cohomological type, by the Lyndon-Hochschild-Serre spectral sequence, it follows that $\Pi[r]\cdot G$ has the same properties. Hence, we conclude by Proposition 2.4. ∎

### v2

@@ 1. Introduction
1.1. The congruence subgroup property
Let $S$ be an oriented connected surface of negative Euler characteristic and finite type and let $G$ be a finite group acting faithfully on $S$. The group $G$ then identifies with a finite subgroup of the mapping class group ${\Gamma}(S)$ of the surface $S$ and let us denote by ${\Gamma}(S)^{G}$ the centralizer of $G$ in ${\Gamma}(S)$. Let $\Pi$ be the fundamental group of $S$ for some choice of base point. There is a natural faithful representation $\rho_{S}\colon\thinspace{\Gamma}(S)\hookrightarrow\operatorname{Out}(\Pi)$ which restricts to a faithful representation $\rho_{S,G}\colon\thinspace{\Gamma}(S)^{G}\hookrightarrow\operatorname{Out}(\Pi)$.
@@ 1.1. The congruence subgroup property
[…]
The congruence subgroup problem for the mapping class group ${\Gamma}(S)$ (resp. for ${\Gamma}(S)^{G}$) asks whether the homomorphism $\hat{\rho}_{S}$ (resp. $\hat{\rho}_{S,G}$) is faithful.
More conventionally, this question can be reformulated by asking whether, for every finite index subgroup $\Gamma$} of ${\Gamma}(S)$ (resp. of ${\Gamma}(S)^{G}$), there exists a finite index characteristic subgroup $K$ of $\Pi$ such that the kernel of the induced representation ${\Gamma}(S)\to\operatorname{Out}(\Pi/K)$ (resp. ${\Gamma}(S)^{G}\to\operatorname{Out}(\Pi/K)$) is contained in $\Gamma$}. The first question is known to have a positive answer for $g(S)\leq 2$ (cf. [1], [2], [6]), where $g(S)$ denotes the genus of $S$.
@@
Let $S$ be an oriented connected hyperbolic surface of finite type and $G$ a finite group acting faithfully on $S$. If the congruence subgroup property holds for the mapping class group ${\Gamma}(S_{/G}^{\circ})$ of the surface $S_{/G}^{\circ}$, then it holds for the centralizer ${\Gamma}(S)^{G}$ of $G$ in the mapping class group ${\Gamma}(S)$ of $S$.
@@
If $g(S_{/G})\leq 2$, then the congruence subgroup property holds for ${\Gamma}(S)^{G}$.
1.2. Conjugacy separability of torsion elements in mapping class groups
An element $x$ of a group $G$ is conjugacy distinguished if, for any $y\in G$, which is not conjugated to $x$ in $G$ (briefly $y\not\sim_{G}x$), there is a finite quotient $\pi\colon\thinspace G\to L$ such that $\pi(y)\not\sim_{L}\pi(x)$.
Similarly, a subgroup $H$ of $G$ is subgroup conjugacy distinguished if, for any subgroup $K$ of $G$, which is not conjugated to $H$ in $G$ (briefly $H\not\sim_{G}K$), there is a finite quotient $\pi\colon\thinspace G\to L$ such that $\pi(H)\not\sim_{L}\pi(K)$.
For a subgroup $H$ of a group $G$, we denote by $Z_{H}(G)$ and $N_{H}(G)$, respectively, the centralizer and the normalizer of $H$ in $G$. For a subset $A$ of a profinite group $R$, we denote by $\overline{A}$ the closure of $A$ for the profinite topology of $R$. We then have (cf. Theorem 5.1 for a slightly more general result):
@@ 1.2. Conjugacy separability of torsion elements in mapping class groups
    The natural monomorphism ${\Gamma}(S)\hookrightarrow{\widehat{\Gamma}}(S)$ induces a bijection between conjugacy classes of solvable finite subgroups. In particular, solvable finite subgroups of ${\Gamma}(S)$ are subgroup conjugacy distinguished.
@@
    Torsion elements of ${\Gamma}(S)$ are conjugacy distinguished.
@@
    For every finite subgroup $G$ of ${\Gamma}(S)$, there holds $Z_{{\widehat{\Gamma}}(S)}(G)=\overline{Z_{{\Gamma}(S)}(G)}$.
@@ 2.1. Normalizers and centralizers of finite subgroups of the mapping class group of a surface
(2) $1\to G\to N_{{\Gamma}(S)}(G)\to{\Gamma}(S_{/G}^{\circ})_{K}\to 1.$
Let $Z(G)$ be the center of the group $G$ and let ${\Gamma}[G]$ be the kernel of the natural representation ${\Gamma}(S_{/G}^{\circ})_{K}\to\operatorname{Out}(G)$. The centralizer ${\Gamma}(S)^{G}\leq N_{{\Gamma}(S)}(G)$ of the subgroup $G$ in the mapping class group ${\Gamma}(S)$ is then described by the short exact sequence:
(3) $1\to Z(G)\to{\Gamma}(S)^{G}\to{\Gamma}[G]\to 1.$
@@
$N_{{\Gamma}(S^{\circ})}(G)\cong N_{{\Gamma}(S)}(G)\hskip 28.45274pt\mbox{and}\hskip 28.45274pt{\Gamma}(S^{\circ})^{G}\cong{\Gamma}(S)^{G}.$
@@ 2.2. The congruence topology
[…]
Let ${\Gamma}(m)^{G}$ be the kernel of the natural representation ${\Gamma}(S)^{G}\to\operatorname{GL}(H_{1}(S,\mathbb{Z}/m))$ and let ${\Gamma}[G,m]$ be its image via the natural epimorphism ${\Gamma}(S)^{G}\to{\Gamma}[G]$. For $m\geq 3$, it is well known that the action of $G$ on $H_{1}(S,\mathbb{Z}/m)$ is faithful. Hence, there holds ${\Gamma}(m)^{G}\cap G=\{1\}$ and so the map ${\Gamma}(S)^{G}\to{\Gamma}[G]$ induces an isomorphism ${\Gamma}(m)^{G}\cong{\Gamma}[G,m]$.
@@
For all $m\geq 2$, the finite index subgroups ${\Gamma}(m)^{G}$ of ${\Gamma}(S)^{G}$ and ${\Gamma}[G,m]$ of ${\Gamma}(S_{/G}^{\circ})$ are both open for the respective congruence topologies.
@@
The statement is obvious for ${\Gamma}(m)^{G}$. For ${\Gamma}[G,m]$, it can be proved with the same argument of [3, Corollary 2.3]. ∎
@@ 2.3. The centralizer of a finite subgroup in the profinite completion of a good group
Let $G$ be a good, residually finite group of finite virtual cohomological type and let $\iota\colon\thinspace G\hookrightarrow\widehat{G}$ be the natural monomorphism. Then, for every finite subgroup $H$ of $G$, we have the identity $Z_{\widehat{G}}(\iota(H))=\overline{\iota(Z_{G}(H))}$.
@@
By [5, (iii) of Theorem A], for all primes $p>0$ and every $p$-torsion element $x$ of $G$, there holds $Z_{\widehat{G}}(\iota(x))=\overline{\iota(Z_{G}(x))}$. The finite group $H$ is generated by a finite subset $\{x_{1},\dots,x_{k}\}$ of $p$-torsion elements (for, possibly, different primes $p>0$). We then have:
$Z_{\widehat{G}}(\iota(H))=\bigcap_{i=1}^{k}Z_{\widehat{G}}(\iota(x_{i}))=\bigcap_{i=1}^{k}\overline{\iota(Z_{G}(x_{i}))}=\overline{\iota(Z_{G}(H))}.$
@@ 3.1. The case of a free action
The congruence subgroup property for ${\Gamma}(S)^{G}$ is equivalent to the statement that the congruence topology on this group, as defined in Section 2.2, is equivalent to the profinite topology. Since, by Lemma 2.2, the subgroup ${\Gamma}(m)^{G}$ of ${\Gamma}(S)^{G}$ is open for the congruence topology, the congruence subgroup property for ${\Gamma}(S)^{G}$ follows if we show that the homomorphism:
[…]
induced by the restriction of $\rho_{S,G}$ to ${\Gamma}(m)^{G}$, is injective for some $m\geq 2$.
In order to prove the above claim, we will need the well known lemma:
@@
[…]
@@
We will also need the following remarks. Since, by hypothesis, the representation $\hat{\rho}_{S_{/G}}\colon\thinspace{\widehat{\Gamma}}(S_{/G})\to\operatorname{Out}({\widehat{\Pi}}_{/G})$ is faithful, its restriction $\hat{\rho}_{S_{/G}}^{(m)}\colon\thinspace{\widehat{\Gamma}}[K,m]\to\operatorname{Out}({\widehat{\Pi}}_{/G})$ is faithful as well for all $m\geq 2$. For $m\geq 3$, the natural isomorphism ${\Gamma}(m)^{G}\cong{\Gamma}[G,m]$ induces an isomorphism of profinite completions ${\widehat{\Gamma}}(m)^{G}\cong{\widehat{\Gamma}}[K,m]$ and $^{\widehat{\Pi}}\cong\widehat{K}$ identifies with a normal finite index subgroup of ${\widehat{\Pi}}_{/G}$ in a way that the representations $\hat{\rho}_{S,G}^{(m)}$ and $\hat{\rho}_{S_{/G}}^{(m)}$ are compatible via the above isomorphisms.
Let then $x\in\ker\hat{\rho}_{S,G}^{(m)}$, put $y:=\hat{\rho}_{S_{/G}}^{(m)}(x)\in\operatorname{Out}({\widehat{\Pi}}_{/G})$ and let $\tilde{y}\in\operatorname{Aut}({\widehat{\Pi}}_{/G})$ be a lift. Since $x\in\ker\hat{\rho}_{S,G}^{(m)}$, by the remarks above, after possibly composing with an inner automorphism of ${\widehat{\Pi}}_{/G}$, we can assume that the restriction of $\tilde{y}$ to $\widehat{K}\equiv{\widehat{\Pi}}$ is the identity. By Lemma 3.1, this implies that $\tilde{y}$ is the identity and so that $y=1$. Since the representation $\hat{\rho}_{S_{/G}}^{(m)}$ is faithful, it follows that $x=1$ as well. This concludes the proof of Theorem A for the case of a free action.
@@ 3.2. The general case
The completions of the groups ${\Gamma}(S)$ and ${\Gamma}(S)^{G}$ with respect to the congruence topologies which we introduced in Section 2.2 are called procongruence completions and denoted by ${\check{\Gamma}}(S)$ and ${\check{\Gamma}}(S)^{G}$, respectively. The group ${\check{\Gamma}}(S)$ is also called the procongruence mapping class group. The congruence subgroup problems for ${\Gamma}(S)$ and ${\Gamma}(S)^{G}$ can then be formulated as the questions of whether the natural epimorphisms ${\widehat{\Gamma}}(S)\to{\check{\Gamma}}(S)$ and ${\widehat{\Gamma}}(S)^{G}\to{\check{\Gamma}}(S)^{G}$ are, respectively, isomorphisms.
By the results of Section 3.1 and with the notations of Section 1, we know that there is a natural isomorphism ${\widehat{\Gamma}}(S^{\circ})^{G}\cong{\check{\Gamma}}(S^{\circ})^{G}$. Therefore, in order to prove the general case of Theorem A, it is enough to show that the natural isomorphism $q^{\circ}\colon\thinspace{\Gamma}(S^{\circ})^{G}\stackrel{{\scriptstyle\sim}}{{\to}}{\Gamma}(S)^{G}$ (cf. Proposition 2.1) induces an isomorphism $\check{q}^{\circ}\colon\thinspace{\check{\Gamma}}(S^{\circ})^{G}\stackrel{{\scriptstyle\sim}}{{\to}}{\check{\Gamma}}(S)^{G}$ between the respective procongruence completions.
The set $R$ of ramification points of the quotient map $p_{G}\colon\thinspace S\to S_{/G}$ determines a point, which we denote by $[R]$, on the unordered configuration space $S[r]$ of $r$ distinct points on the surface $S$, where $r:=\sharp R$.
Let $\Pi[r]:=\pi_{1}(S[r],[R])$. From [6, Corollary 4.7 and Proposition 4.9], it then follows that there is a short exact sequence of profinite groups:
(4) $1\to{\widehat{\Pi}}[r]\to{\check{\Gamma}}(S^{\circ})\stackrel{{\scriptstyle\check{q}^{\circ}}}{{\to}}{\check{\Gamma}}(S)\to 1.$
In order to prove Theorem A, it is then enough to show that the subgroup ${\check{\Gamma}}(S^{\circ})^{G}$ of ${\check{\Gamma}}(S^{\circ})$ has trivial intersection with the image of ${\widehat{\Pi}}[r]$ in ${\check{\Gamma}}(S^{\circ})$. Since ${\check{\Gamma}}(S^{\circ})^{G}$ is contained in the centralizer $Z_{{\check{\Gamma}}(S^{\circ})}(G)$, the conclusion follows if we prove that $Z_{{\check{\Gamma}}(S^{\circ})}(G)\cap{\widehat{\Pi}}[r]=\{1\}$.
@@
Proposition 2.1 implies that ${\Gamma}(S^{\circ})^{G}\cap\Pi[r]=\{1\}$ and then that $Z_{\Pi[r]\cdot G}(G)=Z(G)$. Since $\Pi[r]\cdot G$ contains, as a finite index subgroup, the fundamental group of the ordered configuration space of $n$ points on the surface $S$, which is a good, residually finite group of finite virtual cohomological type, by the Lyndon-Hochschild-Serre spectral sequence, it follows that $\Pi[r]\cdot G$ has the same properties. Hence, we conclude by Proposition 2.4. ∎